### v1

@@ II Magnetic Reconnection in String Wakes
So our model will be based on the dimensions of the string wake that we are considering. In our simulations, we have considered the energy scale for the symmetry-breaking phase transition for Abelian Higgs string to be of the order of $2.2\times 10^{15}$ GeV Lizarraga_2016 . The order of $G\mu$ for such Abelian Higgs cosmic string is therefore approx $10^{-11}-10^{-8}$. There is also the current observational limit from the Planck Collaboration Planck_2016 which gives the upper limit of the value of $G\mu$ as $\sim 10^{-7}$. For the SP model, we can then estimate the length of the reconnection region for the cosmic string case. The total length of the cosmic string wake is taken to be $L$. This is the longest dimension of the wake. Reconnection however does not occur for the full length of the wake. To estimate $l$, we have to obtain the hydromagnetic velocity. This hydromagnetic velocity depends on the density of the plasma and the magnetic field in the wake region. The $l$ of the reconnection region is then obtained from,
$l=\frac{cL}{(C_{H}\sigma L)^{1/2}}.$ (2)
We have considered $c=1$ in all our calculations. Once we know the length of the reconnection region we can obtain the kinetic energy released due to the magnetic reconnection. This energy is what powers the movement of the shock waves in the plasma. So, the kinetic energy will give us the approximate velocties of the shock waves in the plasma.
III Shock-Shock collision in the wake
We initially start with two shocks moving with different velocities away from the point of magnetic reconnection. Since we are interested in the light curves emitted by the GRB, we study the energy evolution with time. We divide the whole process into three stages. In the first stage, the total energy of the plasma is the total energy of the two shock waves moving together. This is equal to the total kinetic energy of the moving shock waves. This kinetic energy however is not constant with time. As the shocks move closer to each other the total energy increases with time. For this we need to know how the velocity of the shock waves are changing with time. As the ejected velocity of the plasma following a magnetic reconnection depends on the magnetic field, we use this knowledge to determine the change in the velocity of the shocks with time. To construct a time dependent model, we write the total energy at a given time in terms of distance and time instead of the velocity. The first shock is assumed to be produced at $t=0$ and the subsequent shock is produced after a period $\Delta t$. The distance at which the collision happens is denoted by $l_{r}$. So both the shocks travel the distance $l_{r}$ from the point at which they were generated at different velocities (and therefore different times $t_{1}$ and $t_{2}$). The kinetic energy of the first shock is given by $\frac{1}{2}\rho_{1}\gamma_{1}\frac{l_{r}^{2}}{t_{1}^{2}}$ and the kinetic energy of the second shock by $\frac{1}{2}\rho_{2}\gamma_{2}\frac{l_{r}^{2}}{t_{2}^{2}}$. Before the collision, the total energy would be the sum of these two energies. At the point of collision, we obtain the maximum energy which is given by the density of the merged shocks $\rho_{m}$, the $\gamma$ factor of the merged shocks $\gamma_{m}$ and the velocity of the shock at the merger point which is again denoted by the distance and the time.
In the numerical model, the time parameter is the variable and we consider different cases for the time difference between the generation of the shock waves. Here, $\rho_{m}=(\rho_{1}+\rho_{2})/2$ and $\gamma_{m}$ is given by,
$\gamma_{m}=\frac{2~{}\sqrt{\rho_{1}^{2}\gamma_{1}^{2}+\rho_{2}^{2}\gamma_{2}^{2}}}{\rho_{1}+\rho_{2}}.$ (3)
Hence to simplify the calculations, we have considered the densities to be integral multiples (e.g, $\rho_{2}=2\rho_{1}$ etc).
Upto the point of collision, the total kinetic energy keeps increasing but after that it will start to decrease as the merged shock gradually dissipates the kinetic energy at collision. The maximum energy occurs at the point of collision. The total energy before the collision depends on time as,
$E_{1}(t)=\frac{c^{2}}{2\sqrt{4\pi\rho}\sigma l_{r}}(\rho_{1}^{2}\Gamma_{1}^{2}+\rho_{2}^{2}\Gamma_{2}^{2})t^{2}.$ (4)
At the collision of the two shocks the maximum energy is given by,
$E_{max}=\frac{\rho_{m}^{2}\Gamma_{m}^{2}c^{2}t_{col}^{2}}{2\sqrt{4\pi\rho}\sigma l_{r}}.$ (5)
@@ III Shock-Shock collision in the wake
Similar to ref.kobayashi we assume once the peak amplitude is obtained, the decrease in the energy is due to the angular spreading of the energy. Assuming the density will remain more or less the same, it will be the velocity which will determine the decrease in energy. As the timestep in the numerical model is small, a Taylor series expansion will result in the approximate decrease of energy with time as $\Delta E\propto(1-\frac{t_{col}^{2}}{t^{2}})$. The constant of proportionality will depend on the same parameters as the maximum energy. So after collision, the energy will decrease as $E(t)-\Delta E$. So the energy expression after the collision is given by,
@@
$E_{2}(t)=\frac{\rho_{m}^{2}\Gamma_{m}^{2}c^{2}t_{col}^{2}}{2\sqrt{4\pi\rho}\sigma l_{r}}\left(1-(1-\frac{t_{col}^{2}}{t^{2}})\right).$ (6)
@@
Fig.[2] shows the resulting pulse that we have obtained.
@@
Figure 2: Change of energy density with time as two shocks waves collide and merge, followed by the decay.
@@
We find that there is a sharp rise followed by a gentle decay. The collision is considered instantaneous and hence obtained the sharp peak. In a more realistic scenario, it would probably happen over a short period of time which will lead to a more rounded peak. For this particular pulse, we have considered the time to be at redshift $z=30$. We have calculated the background density at that temperature and time. We have also taken the average magnetic field to be the background equipartition field at that time. This has a value of $1.95\times 10^{-2}eV^{2}$. The conductivity of the plasma is about $\sigma=0.33cm^{-1}$. The length and width of the cosmic string wake determine the length of the reconnection region. At $z=30$, this turns out to be $l_{r}=1.4\times 10^{10}cm$. Initially, we do our calculations at a fixed redshift, later we check the model by varying the redshift parameter.
IV Multiple collisions and their temporal structure
@@
The important point that we would like to make is that the magnetic reconnection in the cosmic string wake invariably results in mulitple points of magnetic reconnection. This has been illustrated in Fig[1]. So the case of the two shock collision is the simplest possible case. Multiple shocks will be generated and they may be generated from multiple points in the cosmic string wake, so there is the possibility of multiple shock collision in the cosmic string wake. However, it is important to point out that these shocks will have different velocities and therefore different $\gamma$ factors. They may not collide in the same direction. To model this kind of collisions is rather difficult. We have extended our model to include at least a few of these conditions. Extending our two shock model, we have done a simulation where multiple shocks are generated with random values of $\gamma$. These are then made to collide similar to the two shock model that we had before. We can vary the number of shocks colliding one after the other. Though not perfect we have obtained the synthetic light curve for the mulitple collisions also. This is shown in Fig.[3].
@@
Figure 3: Energy density with time as multiple shocks collide and merge, followed by their decay. The shocks have random velocities but generally a faster shock collides with a slower shock in this figure. The number of random collisions is mentioned in the adjoining legends
@@
In Fig.[3], we have shown the light curve for different collision rates. Here, we have assumed that there is no correlation between the multiple shocks and each shock collides only once with the shock closest to it. We have considered all the collisions between a slow shock and a faster shock. As is expected. a larger number of collisions between the shocks only increase the total energy. We also notice that it does not change the order of the total energy generated. So it appears that the number of collisions leading to the GRB will not change the GRB significantly.
@@
We also obtain similar curves with multiple slow moving shocks. By slow moving shocks we mean that the $\gamma$ factor is less than $0.5$c. These are depicted in Fig.[4]. Here too, we get similar light curves, only now the increasing energy is less steep than in the case where one of the shocks was a fast one.
@@
Figure 4: Energy density with time as multiple slow shocks collide and merge, followed by their decay. Here we have kept the shock velocities random but always less than $0.5$c. The number of random collisions is mentioned in the adjoining legends
@@
In all these figures, we have not varied the density of the shocks and considered one of the shocks to have double the density of the other. As mentioned before, this was for the ease of obtaining the resultant $\gamma$ factor of the colliding shocks. As a check, we also did the shock - shock collision for different densities of the shock waves. This is then shown in Fig.[5] with the legends denoting the ratio of the densities of the shocks. As we see in this figure, we are getting different curves for the different ratios with the higher density ratios having higher energy densities.
@@
Figure 5: Energy density with time as multiple shocks of different density ratio collide and merge, followed by their decay. The density ratio is mentioned in the adjoining legends
@@
Finally, we have also checked the effect of varying the background magnetic field. Again this does not change the pattern but changes the overall magnitude and timescale of the GRB. The timescale of the GRB is affected as the emergence of the shocks depends on the magnetic reconnection occuring in the cosmic string wake. For our reconnection model, $\Delta t=\frac{8}{\pi}\frac{l_{r}log(S)}{C_{H}}$. Here, $S$ is the Lundquist number and $C_{H}$ depends on the magnitude of the magnetic field. This is shown in Fig.[6].
@@
Figure 6: Energy density with time as multiple shocks collide and merge for different background magnetic fields. The corresponding background magnetic field is mentioned in the adjoining legends
@@
A general pattern is observed for the synthetic light curve obtained from the shock-shock collision in the magnetized wakes of cosmic strings. The timescale of the collision depends upon various factors. One of the most important factor being the density of the plasma. Since we want to do the calculations for a GRB that is observable, we have chosen $z~{}=~{}30$. However GRB’s are mostly observed at much lower $z$ values. At lower $z$ values, the magnetic reconnection length, the density of the plasma as well as the magnetic field may be different. Keeping the magneitc field value to be the same, we have obtained some graphs for lower values of $z$. This is shown in Fig.[7] for multiple collisions.
@@
Figure 7: Energy density with time as multiple shocks collide and merge for different redshift values. The corresponding values of redshift $z$ are mentioned in the adjoining legends
@@
V Observational possibilities of the light curves

### v2

@@ IV Models of Shock-Shock collisions
Here we are considering that the reconnection occurs in the wake of the cosmic string. So the cosmic string parameter $G\mu$ will determine the length and the width of the reconnection region. We have considered the energy scale for the symmetry-breaking phase transition for Abelian Higgs string to be of the order of $2.2\times 10^{15}$ GeV Lizarraga_2016 . The order of $G\mu$ for such Abelian Higgs cosmic string is therefore approx $10^{-11}-10^{-8}$. There is also the current observational limit from the Planck Collaboration Planck_2016 which gives the upper limit of the value of $G\mu$ as $\sim 10^{-7}$. We can use this to estimate the length of the reconnection region. The total length of the cosmic string wake is taken to be $L$. This is the longest dimension of the wake. Reconnection however does not occur for the full length of the wake. To estimate reconnection length $l$, we have to obtain the Alfven velocity. The Alfven velocity is given by $V_{A}=\frac{B}{\sqrt{\mu\rho}}$, where $\mu$ is the permeability and $\rho$ is the density of the plasma and $B$ is the magnetic field in the wake. Thus, the length of the reconnection region $l$ is obtained as,
$l=\dfrac{cL}{(V_{A}\sigma L)^{1/2}},$ (3)
where, $\sigma$ is the conductivity of the plasma. We will henceforth consider $c=1$ in all our calculations.
To understand how the kinetic energy changes with time, we will use a simple model of shock-shock collision where the mass of the shock changes with time. This change is however small and it occurs due to the continuous streaming of the plasma particles into the wake region. We first obtain the single pulse of energy produced due to the merger of two relativistic shocks with respect to the time interval it is generated in. This time interval is divided into three stages. The first stage is the situation where we have two individual shocks moving towards one another, before the collision, i.e. $t<t_{col}$ and the second stage is when the individual shocks collides and merger happens, i.e. $t=t_{col}$. This is the time at which the kinetic energy reaches it’s peak. Subsequently, in the third stage the energy decreases as the merged shock spreads out, after the collision, i.e. $t>t_{col}$. The energy leads to the acceleration of the particle so the pulse generated in this time interval will determine the luminosity of the GRB emitted due to the radiation from the accelerated particles. One important point that we also tried to incorporate here is the decrease in the magnetic field in this stage. Downstream to the collision of the shocks it is known that the magnetic field decays Zhou_2023 . We have included this decreasing magnetic field in this stage. We now proceed to discuss each stage in some detail.
The first shock is taken to be produced at $t=0$ and the subsequent shock is produced after a time period $\Delta t$. So, the time duration of the collision will be $\Delta t=\dfrac{l_{r}}{2\gamma_{1}^{2}c}$ beniamani . Here, $l_{r}$ denotes the total distance over which the collision takes place and the merger of the shocks happen. So, both the shocks travel the distance $l_{r}$ from the point at which they were generated at different velocities (and therefore different times $t_{1}$ and $t_{2}$). The kinetic energy of the first shock is given by $\dfrac{1}{2}m_{1}\gamma_{1}^{2}c^{2}\frac{B}{(4\pi\rho)^{1/2}\sigma l}$ and the kinetic energy of the second shock by $\dfrac{1}{2}m_{2}\gamma_{2}^{2}c^{2}\frac{B}{(4\pi\rho)^{1/2}\sigma l}$. Before the merger, the total energy would be the sum of these two energies. We are considering only two dimensional shocks here as the shocks are generated from a quasi two-dimensional model. The shocks are generated from the same point and have a radius $l$ at time $t$. As the shocks move with a velocity $v_{s}$, more particles are swept in continuously, so the mass of the shocks will be given on the area $l\times\delta$ by $m_{i}=\rho_{i}\times(l\times\delta)$. The $\delta$ is the width of the reconnection region and is shown in Fig.2. The individual shock velocity $v_{is}^{2}=(l\times\delta)/t^{2}$. The $i$ in the subscript stand for the numbers $1$ or $2$ depending on which shock we are considering. Substituting all these in the kinetic energy equation we get the total energy as a function of time. So in the first stage at $t<t_{col}$, the total energy would be given by
$E_{I}(t)=\dfrac{1}{2}\rho_{1}v_{1s}^{2}\gamma_{1}^{2}c^{2}t^{2}\frac{B}{(4\pi\rho)^{1/2}\sigma l}+\dfrac{1}{2}\rho_{2}v_{2s}^{2}\gamma_{2}^{2}c^{2}t^{2}\frac{B}{(4\pi\rho)^{1/2}\sigma l}.$ (4)
Here, the magnetic field $B$ is considered to be a constant. In the second stage, as the two shocks merge together, the energy reaches the maximum possible value. After this, the density of the combined shock wave is given by the density of the merged shocks $\rho_{m}$, the Gamma factor of the merged shocks $\gamma_{m}$ and the velocity of the shock at the merger point which is again denoted by the distance and the time. So, at the $t=t_{col}$, the distance at which the merger happens is $l_{r}$. Thus, the maximum energy is therefore given by,
$E_{II}(t)=\dfrac{1}{2}\rho_{m}v_{m}^{2}\gamma_{m}^{2}c^{2}t_{col}^{2}\frac{B}{(4\pi\rho)^{1/2}\sigma l_{r}}=E_{max}.$ (5)
Here, $\rho_{m}=(\rho_{1}+\rho_{2})/2$ and $v_{m}\gamma_{m}$ is given by,
$v_{m}\gamma_{m}=\sqrt{\frac{2(\rho_{1}v_{1}^{2}\gamma_{1}^{2}+\rho_{2}v_{2}^{2}\gamma_{2}^{2})}{\rho_{1}+\rho_{2}}}.$ (6)
To simplify the calculations, we have considered the densities of the shocks to be the integral multiples (e.g, $\rho_{2}=2\rho_{1}$ etc). It has been shown that the collision of the shocks will also generate some random magnetic fields which will decay in a short time Zhou_2023 .
Upto the point of collision, the total kinetic energy can increase but after that it will start to decrease as the merged shock gradually dissipates. In several shock collision models, where the particle acceleration has been studied using transport equations, the maximum distribution of the accelerated particles is found to be at the point of collision of the shocks vieu . Since in this case we have not considered the particle distribution function, we denote the total energy of the shocks as the kinetic energy comprising of the mass density and the velocity of the shocks.
@@
Similar to ref.kobayashi , we assume once the peak amplitude is obtained, the decrease in the energy is due to the angular spreading of the shock wave into the background plasma. Assuming the density will remain more or less the same, it will be the velocity which will determine the decrease in energy. As the timestep in the numerical model is small, a Taylor series expansion will result in the approximate increase of energy with time as $\Delta E\propto(1-\frac{t_{col}^{2}}{t^{2}})$. The constant of proportionality will depend on the same parameters as the maximum energy. So after collision, i.e. $t>t_{col}$, the energy will decrease as $E_{max}-\Delta E$. The kinetic energy expression after the collision in the third stage is given by,
$E_{III}(t)=E_{max}\left(1-(1-\frac{t_{col}^{2}}{t^{2}})\right).$ (7)
Usually the kinetic energy leads to the radiation spectra which in most cases is modeled by the Band function. This function consists of a set of exponentially connected broken power laws. Since the magnetic field is an essential part of the kinetic energy dissipated, we would like to model the decrease in the magnetic energy by a time dependent parameter. We consider this decreasing function using a function dependent on time and two other parameters which can be varied. The net result is that the magnetic field should be replaced by a polynomial decaying magnetic field as discussed in ref. Zhou_2023 . Since the expression for $E_{max}$ contains the magnetic field in equation 7. We replace the constant $B$ in this expression with a time dependent magnetic field. The field will decrease with time so at any given time the field will be proportional to $(1-(t_{col}/t))$. Therefore the equation for the energy decay in the third time interval is replaced by
$E_{III}(t)=\dfrac{1}{2}\rho_{m}v_{m}^{2}\gamma_{m}^{2}c^{2}t_{col}^{2}\frac{B(t)}{(4\pi\rho)^{1/2}\sigma l_{r}}\left(1-(1-\frac{t_{col}^{2}}{t^{2}})\right).$ (8)
Here, $B(t)$ denotes the time dependent magnetic field. Since we do not know the actual function, we introduce two variable parameters ($\alpha$ and $\beta$) in the model. We will later on use these parameters to fit the experimental data. The expression therefore becomes
@@
$E_{III}(t)=\dfrac{1}{2}\rho_{m}v_{m}^{2}\gamma_{m}^{2}c^{2}t_{col}^{2}\frac{B}{(4\pi\rho)^{1/2}\sigma l_{r}}\left(1-\left(1-\beta\left({\frac{t_{col}}{t}}\right)^{\alpha}\right)\right).$ (9)
As can be seen, $\alpha=2$ and $\beta=1$ gives us the case of the constant magnetic field. To study the different dependencies we have used these values. The parameters have been used later as free parameters to fit the experimental data. The reason we do this is because we do not know how the magnetic field will actually decay in our present scenario.
@@
Fig.[3] shows the resulting pulse that we have obtained from the above energy expressions for three different stages.
@@
Figure 3: Change of energy fluence with time as two shocks waves collide and merge, followed by the decay.
@@
We find that there is a sharp rise followed by a gentle decay. The collision which leads to the merger is considered instantaneous and hence we have obtained a sharp peak. In a more realistic scenario, it would probably happen over a short period of time which will lead to a more rounded peak. For this particular pulse, we have considered the time to be at redshift $z=6.7$. We have calculated the background plasma density at that temperature and time. We have also taken the average magnetic field to be the background equipartition field at that time. This has a value of $1.95\times 10^{-2}~eV^{2}$. This is the minimum magnetic field at that temperature. Usually, the magnetic field is amplified at least by a factor of $10$ in the wake of the cosmic string. This happens due to the Alfven’s theorem of flux freezing as there is an inward flow of streaming particles in the wake of the cosmic string. An enhanced magnetic field will give us GRB’s in a higher energy band. We have estimated the electrical conductivity of the early universe plasma from the general expression Grasso_2001 ,
$\sigma=\frac{n_{e}e^{2}\tau}{m_{e}},$ (10)
where $n_{e}$ is the electron number density, $e$ is the electron charge, $\tau$ is the mean collision time, and $m_{e}$ is the electron mass. For post-recombination plasma, the electron number density is given by,
$n_{e}(z)\sim 3\times 10^{-10}cm^{-3}~\Omega_{0}h~(1+z)^{3},$ (11)
where $\Omega_{0}$ is the present-time density parameter, $h$ is the Hubble parameter and $z$ is the redshift parameter. The collision time is $\tau\simeq\frac{1}{n_{\gamma}\sigma_{T}}$, where $\sigma_{T}$ Thomson coefficient $\frac{e^{4}}{6\pi m_{e}^{2}}\simeq 6.6\times 10^{-25}cm^{2}$, and $n_{\gamma}=4.2\times 10^{2}cm^{-3}(1+z)^{3}$. Substituting all this in the expression and using $\Omega_{0}=1$ and $h=0.6$, we obtain the electrical conductivity as;
$\sigma=\frac{n_{e}e^{2}}{m_{e}n_{\gamma}\sigma_{T}}\simeq 0.33~{cm}^{-1}.$ (12)
The length and width of the cosmic string wake determine the length of the reconnection region. We have already mentioned that they can be calculated as long as the deficit angle is known. This deficit angle is taken to be $10^{-7}$ from the constraints obtained from the Planck data Planck_2016 . At $z=6.7$, this turns out to be $l=1.4\times 10^{10}$ cm. Initially, we do our calculations at a fixed redshift, later we check the model by varying the redshift parameter.
V Multiple collisions and their temporal structure
@@
The important point that we would like to make is that the magnetic reconnection in the cosmic string wake invariably results in multiple points of magnetic reconnection. This has been illustrated in Fig.[1]. So the case of the two shock collision is the simplest possible case. Multiple shocks will be generated and they may be generated from multiple points in the cosmic string wake, so there is the possibility of multiple shock collision in the cosmic string wake. However, it is important to point out that these shocks will have different velocities and therefore different $\gamma$ factors. They may not collide in the same direction. To model this kind of collisions is rather difficult. We have extended our model to include at least a few of these conditions. Extending our two shock model, we have done a simulation where multiple shocks are generated with random values of $\gamma$. These are then made to collide similar to the two shock model that we had before. We can vary the number of shocks colliding one after the other. Though not perfect, we have obtained the synthetic light curve for the multiple collisions also, as shown in Fig.[4].
@@
Figure 4: Energy fluence with time as multiple shocks collide and merge, followed by their decay. The shocks have random velocities but generally a faster shock collides with a slower shock in this figure. The number of random collisions is mentioned in the adjoining legends.
@@
In Fig.[4], we have shown the light curve for different collision rates. Here, we have assumed that there is no correlation between the multiple shocks and each shock collides only once with the shock closest to it. We have considered all the collisions between a slow shock and a faster shock. As is expected, a larger number of collisions between the shocks only increase the total energy. We also notice that it does not change the order of the total energy generated. So it appears that the number of collisions leading to the GRB will not change the GRBs significantly.
@@
We also obtain similar curves with multiple slow moving shocks. By slow moving shocks we mean that the $\gamma$ factor is less than 0.50.5c. These are depicted in Fig.[5]. Here too, we get similar light curves, only now the increasing energy is less steep than in the case where one of the shocks was a fast one.
@@
Figure 5: Energy fluence with time as multiple slow shocks collide and merge, followed by their decay. Here, we have kept the shock velocities random but always less than 0.50.5c. The number of random collisions is mentioned in the adjoining legends
@@
As mentioned before, the background densities of these shock waves have been taken as integral multiples of each other. This was for the ease of obtaining the resultant $\gamma$ factor of the colliding shocks. As a check, we also did the shock-shock collision for different densities of the shock waves. This is then shown in Fig.[6] with the legends denoting the ratio of the densities of the shocks. As we see in this figure, we are getting different curves for the different ratios with the higher density ratios having higher energy fluence.
@@
Figure 6: Energy fluence with time as multiple shocks of different density ratio collide and merge, followed by their decay. The density ratio is mentioned in the adjoining legends
@@
Finally, we have also checked the effect of varying the background magnetic field. Again this does not change the pattern but changes the overall magnitude and timescale of the GRB. The timescale of the GRB is affected as the emergence of the shocks depends on the magnetic reconnection occurring in the cosmic string wake. For our fast reconnection model, $\Delta t=\frac{8}{\pi}\frac{l_{r}log(S)}{V_{A}}$. Here, $S$ is the Lundquist number and $V_{A}$ depends on the magnitude of the magnetic field and plasma density. This is shown in Fig.[7].
@@
Figure 7: Energy fluence with time as multiple shocks collide and merge for different background magnetic fields. The corresponding background magnetic field is mentioned in the adjoining legends
@@
A general pattern is observed for the synthetic light curve obtained from the shock-shock collision in the magnetized wakes of cosmic strings. The timescale of the collision depends upon various factors. One of the most important factor being the density of the plasma. Since we want to do the calculations for a GRB that is observable, we have chosen redshift $z=6.7$. However GRB’s are mostly observed at much lower $z$ values. At different redshift $z$ values, the magnetic reconnection length, the density of the plasma as well as the magnetic field may be different. Keeping the magnetic field value to be the same, we have obtained some graphs for corresponding redshift values of $z$. This is shown in Fig.[8] for multiple collisions.
@@
Figure 8: Energy fluence with time as multiple shocks collide and merge for different redshift values. The corresponding values of redshift $z$ are mentioned in the adjoining legends
@@
VI Observational possibilities of the light curves